OpenStudy (anonymous):

use logarithmic differentiation to do: y=(x^2 +1/x^2 -1)^1/4

OpenStudy (anonymous):

\[y=(\frac{ x ^{2}+1 }{ x^{2}-1 })^{\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

\[\ln(y)=\frac{1}{4}\left(\ln(x^2+1)-\ln(x^2-1)\right)\] is a start

OpenStudy (anonymous):

could you walk me through it?

OpenStudy (anonymous):

then take the derivative piece by piece leave the \(\frac{1}{4}\) out front

OpenStudy (anonymous):

the derivative of \(\ln(f(x))\) is \(\frac{f'(x)}{f(x)}\) so the next step would be to take the derivative of \(\ln(y)\) and get \[\frac{1}{4}\left(\frac{2x}{x^2+1}-\frac{2x}{x^2-1}\right)\]

OpenStudy (anonymous):

then to finish, multiply by the original function

OpenStudy (anonymous):

thanks